OpenStudy (anonymous):

FAN AND MEDAL: Which statement is true about the sum square root of 50 + square root of 6? It is rational and equal to 2. It is rational and equal to 8. It is irrational and equal to 2square root of 2. It is irrational and equal to 8square root of 2.

OpenStudy (anonymous):

@esshotwired @phi @undeadknight26

OpenStudy (anonymous):

@thomaster @UnkleRhaukus @sweetburger Can any of y'all help me?

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

@satellite73 @ganeshie8

OpenStudy (anonymous):

i think its A

OpenStudy (anonymous):

\[\sqrt{50}+\sqrt6\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

it is none of your choices

OpenStudy (anonymous):

it is irrational

OpenStudy (anonymous):

and you can write \[\sqrt{50}=5\sqrt2\]add to get \[5\sqrt2+\sqrt6\]

OpenStudy (anonymous):

perhaps it is \[\sqrt{50}+\sqrt{18}\]

OpenStudy (anonymous):

that would be \[5\sqrt2+3\sqrt2=8\sqrt2\]

OpenStudy (anonymous):

Okay..I will definitely try that. Thanks for the help!!!

OpenStudy (bradely):

square root 50 = square root (25*2) =square root (25)*square root (2) =5squar root 2 square root 50 +square root 6 = 5squar root 2+ square root 6